# Achieving Resolution-Agnostic DNN-based Image Watermarking: A Novel Perspective of Implicit Neural Representation

### v1

@@ 1. Introduction
RAIMark comprises three key stages, as shown in Figure 3. In stage 1, we generate the implicit neural representation by fitting it with an arbitrary-resolution image. Stage 2 involves pre-training a decoder that is independent of the image and capable of extracting watermarks from images of any resolution. In stage 3, we first generate a pre-defined message. Then, we embed the watermarks into the INR by fine-tuning the model using the pre-trained decoder to ensure the same message can be obtained from images sampled by the same INR. We obtain watermarked images of different resolutions during testing by feeding different parameters to the sampler. Our contributions are summarized as follows:
@@ 4.2. Pre-training the Watermark Decoder
(5) $\mathcal{L}=-\sum_{i=1}^{n}M_{i}\cdot\log(\sigma(m_{i}))+(1-M_{i})\cdot\log(1-\sigma(m_{i})),$
@@ 4.3. Fine-tuning the Implicit Neural Representation
First, we generate a pre-defined message $m=(m_{1},\dots,m_{n})\in\{0,1\}^{n}$ for a given INR $F_{im}$. We need to save this message and use the same message during validation and testing. Then, we feed the fine-tuning INR $F_{wm}$ to a sampler $\mathcal{S}_{(H,W)}$ that outputs an image $I_{s}\in\mathbb{R}^{3\times H\times W}$. Moreover, $I_{s}$ is the image with watermarks. During training, we change the resolution of $H$ and $W$ and add samples of different resolutions to improve generalization. The noise layers distort the sampled image $I_{n}=\mathcal{N}(I_{s})$, and the pre-trained decoder extracts a message $m^{\prime}=\mathcal{D}(I_{n})$. The loss function of the message part is the BCE between extracted message $m^{\prime}$ and pre-defined message $m$:
@@ 4.4. Noise Layers
Since watermarked images tend to suffer from various distortions in real-life scenarios, we added a noise layer in the training process to enhance the robustness of our model. The details of all noises we used in our method are shown in Table 1. We classify the noise into differentiable or non-differentiable depending on its realization. Differentiable noise means that after applying themselves to watermarked images, we can typically perform the reverse process when training. Moreover, with non-differentiable noises like JPEG compression, the backward propagation fails to produce the corresponding gradient because of the saving and reading of the images. Here, we choose Forward ASL (Zhang et al., 2021), which is compatible with non-differentiable noises. Forward ASL first calculates the difference between noised image $I_{n}$ and watermarked image $I_{w}$, $I_{diff}=I_{n}-I_{w}$. Here, $I_{diff}$ has no gradient. The new noised image $I_{n}^{\prime}=I_{w}+I_{diff}$ is the input of the decoder. The non-differentiable noise does not participate in the gradient propagation during the backward process. Therefore, the gradient can be back-propagated through the noise layer.
@@ 5.1. Implementation Details
In the fine-tuning process, to ensure invisibility and robustness over different resolutions, we fine-tune the INR in three samples, $256\times 256$, $384\times 384$, and $512\times 512$. We utilize Adam optimizer with a learning rate of $5\times 10^{-5}$. We fine-tune the INR under non-geometric attacks and geometric attacks. Our choice for the coefficients $\lambda_{msg}$ and $\lambda_{img}$ are $5\times 10^{5}$ and $3\times 10^{3}$. We fine-tune 500 epochs and choose the best-performing model as the watermarked INR $F_{wm}$.
@@ 5.4.2. Combined Distortions
We observe that previous works are vulnerable to at least one image attack. For example, DWSF is vulnerable to cropping attack because its bit accuracy drops to 51.33 under cropping attack. The reason is that DWSF only hides watermarks within selected blocks. Once selected blocks are cropped, the watermark cannot be verified. MBRS is also vulnerable to cropping attack because a specific region of the watermarked image only hides part of the binary message. Once the region is cropped, the part of the binary message it hides cannot be recovered from other image regions.
@@ 5.5.2. Results on invisibility and Robustness
In this section, we analyze the invisibility and robustness of our model on six different samples: $128\times 128$, $256\times 256$, $512\times 512$, $480\times 854$ (480p), $720\times 1280$ (720p) and $1080\times 1920$ (1080p).

### v2

@@ 1. Introduction
RAIMark comprises three key stages, as shown in Figure 3. In stage 1, we generate the implicit neural representation by fitting it with an arbitrary-resolution image. Stage 2 involves pre-training a decoder that is independent of the image and capable of extracting watermarks from images of any resolution. In stage 3, we first generate a watermark message. Then, we embed the watermarks into the INR by fine-tuning the model using the pre-trained decoder to ensure the same message can be obtained from images sampled by the same INR. We obtain watermarked images of different resolutions during testing by feeding different parameters to the sampler. Our contributions are summarized as follows:
@@ 4.2. Pre-training the Watermark Decoder
(5) $\mathcal{L}=-\frac{1}{n}\sum_{i=1}^{n}M_{i}\cdot\log(\sigma(m_{i}))+(1-M_{i})\cdot\log(1-\sigma(m_{i})),$
@@ 4.3. Fine-tuning the Implicit Neural Representation
First, we generate a message $m=(m_{1},\dots,m_{n})\in\{0,1\}^{n}$ for a given INR $F_{im}$. We need to save this message and use the same message during validation and testing. Then, we feed the fine-tuning INR $F_{wm}$ to a sampler $\mathcal{S}_{(H,W)}$ that outputs an image $I_{s}\in\mathbb{R}^{3\times H\times W}$. Moreover, $I_{s}$ is the image with watermarks. During training, we change the resolution of $H$ and $W$ and add samples of different resolutions to improve generalization. The noise layers distort the sampled image $I_{n}=\mathcal{N}(I_{s})$, and the pre-trained decoder extracts a message $m^{\prime}=\mathcal{D}(I_{n})$. The loss function of the message part is the BCE between extracted message $m^{\prime}$ and generated message $m$:
@@ 4.4. Noise Layers
Since watermarked images tend to suffer from various distortions in real-life scenarios, we added a noise layer in the training process to enhance the robustness of our model. The details of all noises we used in our method are shown in Table 1. We classify the noise into differentiable or non-differentiable depending on its realization. Differentiable noise means that after applying themselves to watermarked images, we can typically perform the reverse process when training. Moreover, with non-differentiable noises like JPEG compression, the backward propagation fails to produce the corresponding gradient because of the saving and reading of the images. Here, we choose Forward ASL (Zhang et al., 2021), which is compatible with non-differentiable noises. Forward ASL first calculates the difference between noised image $I_{n}$ and watermarked image $I_{w}$, $I_{diff}=I_{n}-I_{w}$. Here, $I_{diff}$ does not participate in gradient calculation. The new noised image $I_{n}^{\prime}=I_{w}+I_{diff}$ is the input of the decoder. The non-differentiable noise does not participate in the gradient propagation during the backward process. Therefore, the gradient can be back-propagated through the noise layer.
@@ 5.1. Implementation Details
In the fine-tuning process, to ensure invisibility and robustness over different resolutions, we fine-tune the INR in three samples, $256\times 256$, $384\times 384$, and $512\times 512$. We utilize Adam optimizer with a learning rate of $5\times 10^{-5}$. We fine-tune the INR under non-geometric attacks and geometric attacks. Our choice for the coefficients $\lambda_{img}$ and $\lambda_{msg}$ are $5\times 10^{5}$ and $3\times 10^{3}$. We fine-tune 500 epochs and choose the best-performing model as the watermarked INR $F_{wm}$.
@@ 5.4.2. Combined Distortions
We observe that previous works are vulnerable to at least one image attack. For example, DWSF is vulnerable to cropping attack because its bit accuracy drops to 51.33% under cropping attack. The reason is that DWSF only hides watermarks within selected blocks. Once selected blocks are cropped, the watermark cannot be verified. MBRS is also vulnerable to cropping attack because a specific region of the watermarked image only hides part of the binary message. Once the region is cropped, the part of the binary message it hides cannot be recovered from other image regions.
@@ 5.5.2. Results on invisibility and Robustness
In this section, we analyze the invisibility and robustness of our model on six different samples: $256\times 256$, $384\times 384$, $512\times 512$, $480\times 854$ (480p), $720\times 1280$ (720p) and $1080\times 1920$ (1080p).